OpenStudy (anonymous):

x/3= 2/x what is the value of x? I'm looking more for the process than the answer (although that would be nice too :)

OpenStudy (anonymous):

\[x^2=6\]

OpenStudy (turingtest):

multiply both sides by 3 x=6/x then by x \[x^2=6\]\[x=\sqrt{6}\]

OpenStudy (turingtest):

a faster way is to see ahead that by multiplying the whole thing by 3x gets rid of all the fractions at once, but that takes a little practice to see so quickly.

OpenStudy (anonymous):

Thank you, both answers helpful :)